The cost of 2 bagels and 2 cans of juice is $4.40. The cost of 3 bagels and 4 cans of juice is $7.80
a. write a system f equations to represent this situation.
b. Solve the system of equations by substitution. Explain what the solution means.

Answer by FrankM(1040)

2B+2J=440
3B+4J=780
Double both sides of first equation;
4B+4J=880
3B+4J=780 (and subtract the second equation from the first here)
B=100 A bagel is $1.00
2(100)+2J=440
200+2J=440
2J=240
J=120 Juice is $1.20